# Durian: Dual Reference-guided Portrait Animation with Attribute Transferthanks: Project page: https://hyunsoocha.github.io/durian

### v1

@@ Abstract
We present Durian, the first method for generating portrait animation videos with facial attribute transfer from a given reference image to a target portrait in a zero-shot manner. To enable high-fidelity and spatially consistent attribute transfer across frames, we introduce dual reference networks that inject spatial features from both the portrait and attribute images into the denoising process of a diffusion model. We train the model using a self-reconstruction formulation, where two frames are sampled from the same portrait video: one is treated as the attribute reference and the other as the target portrait, and the remaining frames are reconstructed conditioned on these inputs and their corresponding masks. To support the transfer of attributes with varying spatial extent, we propose a mask expansion strategy using keypoint-conditioned image generation for training. In addition, we further augment the attribute and portrait images with spatial and appearance-level transformations to improve robustness to positional misalignment between them. These strategies allow the model to effectively generalize across diverse attributes and in-the-wild reference combinations, despite being trained without explicit triplet supervision. Durian achieves state-of-the-art performance on portrait animation with attribute transfer, and notably, its dual reference design enables multi-attribute composition in a single generation pass without additional training.
@@ 1 Introduction
Personalized appearance editing, such as trying on glasses or exploring new hairstyles, is increasingly common in AR applications. However, most existing tools rely on static previews or simplified 2D templates, limiting realism and expressiveness. A single image is often insufficient for users to judge whether a new look truly suits them, particularly for more significant changes like facial hair or hairstyles. Users naturally want to see how a style behaves with facial expressions or looks from different angles. This motivates the need for a method that can animate a single portrait image using the appearance of a reference face with the desired attributes.
Figure 1: Overview of Durian. Given a portrait and a reference image specifying target attributes (e.g., hairstyle, eyeglasses), our Durian generates an animatable 2D portrait video with composable visual attributes in a single-stage pipeline.
Recent diffusion-based methods have enabled high-quality image editing by injecting visual attributes from one image into another. Exemplar-based inpainting approaches (Yang et al., 2023; Chen et al., 2024) support applications like hairstyle transfer and virtual try-on. However, these methods are designed for static image editing and often require manual input (e.g., masks or user scribbles) to localize the transfer region. Extensions to video (Fang et al., 2024; Tu et al., 2025) still assume that the target regions are easy to specify, which does not hold for dynamic and deformable facial attributes. Other methods (Zhang et al., 2025; Chung et al., 2025) focus specifically on hairstyle transfer by constructing task-specific training data. For example,  Zhang et al. (2025) synthesizes triplets by predicting a bald version of a portrait and generating reference hair images via a pretrained diffusion model, but this process is not scalable and hard to extend to other attributes. Moreover, these methods are confined to static images and are unable to generate dynamic expression or pose-aware animations.
In this work, we propose a zero-shot framework for portrait animation with cross-identity attribute transfer. Given only a single portrait of the source identity and another image depicting the desired attribute (e.g., hairstyle, beard, or eyeglasses), our method generates a video of the source face turning naturally while seamlessly adopting the target attribute, as shown in figure 1. A straightforward solution might involve constructing triplet data—consisting of a reference attribute image, a target identity before the edit, and the desired edited version—but building such triplets for a wide variety of facial attributes lacks scalability. The problem becomes even more challenging for video, where consistency across frames is hard to guarantee. Alternatively, attribute transfer can be formulated as an inpainting or object insertion task. However, this requires accurate region masks, which are difficult to define in a general and scalable way, as the spatial extent of the attribute (e.g., long vs. short hair) can vary significantly. These challenges call for a flexible and generalizable framework that performs attribute-guided animation without requiring triplet training data or frame-level masks.
To address these challenges, we extend self-reference training setups commonly used in exemplar-based inpainting (Yang et al., 2023) to the video domain for portrait animation with attribute transfer. Leveraging in-the-wild portrait videos, we randomly sample one frame and decompose it into two inputs: an attribute-only image and an attribute-masked portrait image. These are fed into a diffusion model, which is trained to generate the remaining frames of the video. Notably, the attribute mask is required only for the two input images—not for the output frames—allowing the model to implicitly propagate the desired attribute throughout the animation. To accommodate attribute transfer with varying spatial extent, we expand the mask region in the attribute-masked portrait image, exposing the model to a wider range of insertion scenarios. To further improve robustness against the pose and alignment variations commonly seen in in-the-wild pairs, we introduce a tailored augmentation strategy that perturbs the spatial layout of the input pair. This allows the model to generalize across diverse attribute-portrait configurations. Our setup enables robust, zero-shot animation with diverse attributes and minimal annotation effort.
Our contributions are summarized below:
  • We propose the first method to directly generate portrait animation video with transferred attributes from two images, generalized over diverse facial attributes, not limited to hair.
  • We design a self-reconstruction training for attribute transfer that enables scalable learning on in-the-wild data, enhanced by spatial mask expansion and tailored augmentation to improve robustness against pose and attribute variations.
  • Our framework naturally supports multi-attribute composition within a single generation pass, without requiring any additional training or fine-tuning.
@@ 2 Related Work
Face Image and Video Editing.
@@
Diffusion-based Virtual Try-On.
Diffusion-based try-on methods typically formulate editing as masked inpainting, where reference content is inserted into a target image using explicit masks (Yang et al., 2023; Chen et al., 2024; Mou et al., 2025; Chen et al., 2025; Song et al., 2025). These approaches have been adapted to domain-specific tasks such as clothing (Kim et al., 2024; Li et al., 2024; Chong et al., 2024), hair (Zhang et al., 2025; Chung et al., 2025), and makeup (Zhang et al., 2024b). While effective for static images, they rely on category labels or mask annotations. Video extensions (Fang et al., 2024; Tu et al., 2025) apply per-frame inpainting with post-hoc smoothing, but predefined masks are hard to specify for deformable facial attributes that vary over time. In contrast, our model performs attribute transfer and animation jointly in a single forward pass, conditioned only on a pair of reference images and a facial keypoint sequence. This eliminates the need for per-frame masks, text prompts, or category labels, enabling zero-shot transfer of diverse facial attributes.
@@
Portrait animation aims to generate motion from a static image, typically guided by facial keypoints, audio, or motion trajectories. Early methods rely on GANs with implicit keypoint modeling (Guo et al., 2024; Wang et al., 2021), while recent approaches use diffusion models (Hu, 2024; Zhu et al., 2024; Yang et al., 2025) for improved realism and temporal stability. These methods primarily focus on reenactment and appearance preservation. Others incorporate paired motion (Xie et al., 2024) or audio (Yang et al., 2025), but require multi-stage inference or fine-tuning. Our model unifies facial attribute transfer and motion generation in a single framework. Without subject-specific tuning, it synthesizes identity-preserving, photorealistic videos from diverse attribute references and keypoint-driven motion, enabling compositional edits and smooth animation in a single pass.
@@ 3 Method
Figure 2: Overview of Training Pipeline. Given an attribute-masked portrait image $\tilde{\mathbf{I}}_{\mathrm{port}}$ and an attribute-only image $\tilde{\mathbf{I}}_{\mathrm{attr}}$, our Durian synthesizes a portrait animation with the transferred attribute. These inputs are constructed by randomly sampling two frames from a training video and applying the estimated masks. A sequence of facial keypoints ${\bm{k}_{\tau}}_{\tau=1}^{F}$ is extracted from the video to guide the motion of the generated animation. During generation, spatial features from PRNet and ARNet are fused via spatial attention into the DNet, enabling identity-preserving and attribute-consistent video synthesis.
3.1 Overview: Learning Attribute Transfer from Self-Recon.
We propose a diffusion-based generative framework for portrait animation with cross-identity attribute transfer. At a high level, the model generates an $F$-frame animation sequence $\mathbf{V}=\{\mathbf{I}_{\tau}\}_{\tau=1}^{F}$ as:
@@
conditioned on an attribute image $\mathbf{I}_{\mathrm{attr}}$, a portrait image $\mathbf{I}_{\mathrm{port}}$, and a sequence of driving facial keypoint images $\mathbf{K}=\{\bm{k}_{\tau}\}_{\tau=1}^{F}$. Each reference image is accompanied by a binary mask: $\mathbf{M}_{\mathrm{attr}}$ localizes the attribute region (e.g., hair or glasses) in the reference image, while $\mathbf{M}_{\mathrm{port}}$ specifies the candidate region in the portrait where the attribute will be transferred. Using these masks, we construct two masked inputs: the attribute-only image $\tilde{\mathbf{I}}_{\mathrm{attr}}=\mathbf{I}_{\mathrm{attr}}\odot\mathbf{M}_{\mathrm{attr}}$, where only the attribute region is preserved, and the attribute-masked portrait image $\tilde{\mathbf{I}}_{\mathrm{port}}=\mathbf{I}_{\mathrm{port}}\odot(1-\mathbf{M}_{\mathrm{port}})$, where the corresponding region is removed. These masked inputs are fed into the Dual ReferenceNet, consisting of the Attribute ReferenceNet (ARNet) and Portrait ReferenceNet (PRNet), which extract multi-scale spatial features. These features are then injected into a diffusion-based generator, the Denoising UNet (DNet), to synthesize the remaining frames of the video, guided by the sequence of keypoints (Section˜3.2).
@@
To enable training without requiring explicitly annotated triplets (i.e., combinations of a target attribute image, an original portrait image, and an edited portrait image), we adopt a self-reconstruction strategy based on portrait videos (Yu et al., 2023; Xie et al., 2022). Specifically, we simulate attribute transfer by sampling two frames $\mathbf{I}_{\mathrm{attr}}$ and $\mathbf{I}_{\mathrm{port}}$ from the same video, treating one as the attribute reference and the other as the target portrait. We then construct the masked inputs $\tilde{\mathbf{I}}_{\mathrm{attr}}$ and $\tilde{\mathbf{I}}_{\mathrm{port}}$ using the same masking formulation as in inference, based on a segmentation mask of a randomly selected attribute. Although the two frames originate from the same identity, this complementary masking enforces role separation between identity and attribute inputs and encourages the model to learn meaningful mappings from attribute and identity features to output frames without cross-identity supervision. To enhance the model’s ability to generalize beyond the self-attribute transfer setup, we introduce an augmentation pipeline that improves robustness to variations in attribute shape, position, and appearance (Section˜3.3).
@@
At inference time, we estimate aligned attribute masks using a 3D animatable avatar to mitigate spatial misalignment between the attribute image and the portrait. Conditioned on the masked reference images and the driving keypoint sequence, our model synthesizes temporally coherent portrait animations with attribute transfer. Notably, our design supports multi-attribute composition and interpolation between attributes in a single generation pass, without requiring additional training or post-processing (Section˜3.4). Fig.˜1 shows our generated portrait animations with attribute transfer.
@@ 3.2 Model Architecture: Dual ReferenceNet
Inspired by recent approaches (Guo et al., 2023; Hu, 2024; Zhu et al., 2024) that leverage ReferenceNet to inject spatial features into diffusion models, we present a Dual ReferenceNet architecture tailored for portrait animation with attribute transfer. Unlike previous work, our model includes two separate encoders: Attribute ReferenceNet (ARNet) and Portrait ReferenceNet (PRNet), each sharing the same architecture as the Denoising U-Net (DNet) in the diffusion model. The networks are adapted from the U-Net (Long et al., 2015) used in latent diffusion models (Rombach et al., 2022), where each block consists of convolutional layers, followed by self-attention and cross-attention modules. The overall architecture is illustrated in Fig.˜2.
@@
Given an attribute image $\mathbf{I}_{\mathrm{attr}}\in\mathbb{R}^{3\times H\times W}$ and a portrait image $\mathbf{I}_{\mathrm{port}}\in\mathbb{R}^{3\times H\times W}$, along with their binary masks $\mathbf{M}_{\mathrm{attr}}\in\mathbb{R}^{1\times H\times W}$ and $\mathbf{M}_{\mathrm{port}}\in\mathbb{R}^{1\times H\times W}$, which localize the attribute region and the candidate transfer region respectively, we construct two masked inputs: the attribute-only image $\tilde{\mathbf{I}}_{\mathrm{attr}}=\mathbf{I}_{\mathrm{attr}}\odot\mathbf{M}_{\mathrm{attr}}$, where only the attribute region is preserved, and the attribute-masked portrait image $\tilde{\mathbf{I}}_{\mathrm{port}}=\mathbf{I}_{\mathrm{port}}\odot(1-\mathbf{M}_{\mathrm{port}})$, where the corresponding candidate region is removed. We then encode these masked images into latent representations using the pretrained VAE from the latent diffusion model (Rombach et al., 2022), yielding $\bm{z}_{\mathrm{attr}},\bm{z}_{\mathrm{port}}\in\mathbb{R}^{c\times h\times w}$. The corresponding masks $\mathbf{M}_{\mathrm{attr}},\mathbf{M}_{\mathrm{port}}$ are downsampled to match the latent resolution, producing $\bm{m}_{\mathrm{attr}},\bm{m}_{\mathrm{port}}\in\mathbb{R}^{1\times h\times w}$. These downsampled masks are concatenated with the latents along the channel dimension to form $(c+1)$-channel inputs:
$\tilde{\bm{z}}_{\mathrm{attr}}=\mathrm{concat}_{\mathrm{c}}(\bm{z}_{\mathrm{attr}},\bm{m}_{\mathrm{attr}}),\quad\tilde{\bm{z}}_{\mathrm{port}}=\mathrm{concat}_{\mathrm{c}}(\bm{z}_{\mathrm{port}},\bm{m}_{\mathrm{port}}),$ (2)
where $\tilde{\bm{z}}_{\mathrm{attr}},\tilde{\bm{z}}_{\mathrm{port}}\in\mathbb{R}^{(c+1)\times h\times w}$.
@@
The augmented latents are passed to ARNet $\mathcal{E}_{\mathrm{attr}}$ and PRNet $\mathcal{E}_{\mathrm{port}}$, with the same architecture as DNet, to extract multi-scale feature maps after convolutional layers of each block:
$\mathcal{F}_{\mathrm{attr}}$ $\coloneq\{\mathbf{F}_{\mathrm{attr}}^{l}\}_{l=1}^{L}=\mathcal{E}_{\mathrm{attr}}(\tilde{z}_{\mathrm{attr}};\Theta_{\mathrm{attr}}),$ (3)
$\mathcal{F}_{\mathrm{port}}$ $\coloneq\{\mathbf{F}_{\mathrm{port}}^{l}\}_{l=1}^{L}=\mathcal{E}_{\mathrm{port}}(\tilde{z}_{\mathrm{port}};\Theta_{\mathrm{port}}),$ (4)
where $\Theta_{\{\mathrm{attr,port}\}}$ are the parameters of Dual ReferenceNet.
Let $\mathbf{F}_{t}^{\tau,l}\in\mathbb{R}^{c_{l}\times h_{l}\times w_{l}}$ denote the feature map of the frame $\tau$ at the $l$-th block of the denoising U-Net. While the original denoising U-Net includes a self-attention layer at each resolution, we replace it with our spatial attention to integrate identity and attribute features in a spatially-aware manner. We denote width-wise concatenation as $\mathrm{concat}_{\mathrm{w}}(\cdot)$, and define our spatial attention $\mathrm{SA}(\cdot,\cdot,\cdot)$ as:
@@
$\bar{\mathbf{F}}_{t}^{l}=\mathrm{SA}(\mathbf{F}_{t}^{\tau,l},\mathbf{F}_{\mathrm{port}}^{l},\mathbf{F}_{\mathrm{attr}}^{l})=\mathrm{Attention}(\mathbf{W}_{Q}\mathbf{F}_{t}^{\tau,l},\mathbf{W}_{K}\mathbf{F}_{\mathrm{ref},t}^{\tau,l},\mathbf{W}_{V}\mathbf{F}_{\mathrm{ref},t}^{\tau,l}),$ (6)
where $\bar{\mathbf{F}}_{t}^{\tau,l}\in\mathbb{R}^{c_{l}\times h_{l}\times w_{l}}$ is the feature map after the spatial attention,
$\mathrm{Attention}(Q,K,V)=\mathrm{softmax}({QK^{\top}}/{\sqrt{d}})V$ (7)
is the standard scaled dot-product attention (Vaswani et al., 2017), $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ are linear projection layers. This width-wise concatenation preserves spatial resolution and allows the model to attend across all positions in the combined reference and target features. As a result, the model can leverage both attribute and portrait guidance at every generation step, enabling fine-grained and semantically consistent attribute transfer across identities.
@@
After applying spatial attention, we further inject semantic guidance into both the Dual ReferenceNet and the denoising U-Net via cross-attention. For ARNet, we use the CLIP (Radford et al., 2021) embedding of the attribute-only image $\tilde{\mathbf{I}}_{\mathrm{attr}}$ as the attribute embedding $\bm{\phi}_{\mathrm{attr}}$, which is injected via cross-attention into each block of ARNet. For PRNetand DNet, we follow StableAnimator (Tu et al., 2024) and construct a portrait embedding $\bm{\phi}_{\mathrm{port}}$ by combining ArcFace (Deng et al., 2019) and CLIP embeddings of the attribute-masked portrait image $\tilde{\mathbf{I}}_{\mathrm{port}}$. This embedding is injected into both PRNet and DNet to enhance identity preservation. We define the cross-attention operation $\mathrm{CA}(\cdot,\cdot)$ as:

### v2

# Durian: Dual Reference Image-Guided Portrait Animation with Attribute Transfer

@@ Abstract
We present Durian, the first method for generating portrait animation videos with cross-identity attribute transfer from one or more reference images to a target portrait. Training such models typically requires attribute pairs of the same individual, which are rarely available at scale. To address this challenge, we propose a self-reconstruction formulation that leverages ordinary portrait videos to learn attribute transfer without explicit paired data. Two frames from the same video act as a pseudo pair: one serves as an attribute reference and the other as an identity reference. To enable this self-reconstruction training, we introduce a Dual ReferenceNet that processes the two references separately and then fuses their features via spatial attention within a diffusion model. To make sure each reference functions as a specialized stream for either identity or attribute information, we apply complementary masking to the reference images. Together, these two components guide the model to reconstruct the original video, naturally learning cross-identity attribute transfer. To bridge the gap between self-reconstruction training and cross-identity inference, we introduce a mask expansion strategy and augmentation schemes, enabling robust transfer of attributes with varying spatial extent and misalignment. Durian achieves state-of-the-art performance on portrait animation with attribute transfer. Moreover, its dual reference design uniquely supports multi-attribute composition and smooth attribute interpolation within a single generation pass, enabling highly flexible and controllable synthesis. Project page: https://hyunsoocha.github.io/durian
Figure 1: Portrait Animation with Attribute Transfer. Given a portrait image and single or multiple reference images specifying target attributes (e.g., hairstyle, eyeglasses), our method generates a portrait animation with facial attribute transfer conditioned on a keypoint sequence.
@@ 1 Introduction
Personalized appearance editing, such as virtually trying on glasses or experimenting with new hairstyles, is becoming a key feature of virtual styling applications. However, most existing solutions are highly specialized and limited in scope. Hairstyle preview apps typically rely on fixed templates, which may look realistic from a single view but fail to adapt to head pose or expression changes. Glasses try-on systems often depend on pre-scanned 3D product models, restricting users to a predefined catalog. Furthermore, these systems focus on a single attribute and cannot combine multiple elements, such as hair, glasses, or hats, within a unified experience.
A key challenge in building such a system is obtaining suitable training data. Disentangling identity from attributes ideally requires paired images of the same person with different attributes, which are rarely available and expensive to collect at scale. This difficulty grows exponentially for multiple attributes, as capturing all combinations quickly becomes infeasible. For example, Li et al. (2023) collects multi-view images of subjects wearing different eyeglasses to model realistic glasses try-on, but the dataset remains too limited to generalize broadly. Zhang et al. (2025) propose a synthetic pipeline that predicts a bald version of a portrait and generates reference hair images using a pretrained diffusion model. However, this approach is not easily scalable beyond hair.
This naturally raises the question: can we train a model for portrait animation with attribute transfer without any explicit attribute-paired data? Motivated by this question, we propose a self-reconstruction framework that learns this task directly from widely available in-the-wild portrait videos. During training, we randomly sample two frames from a single video: one as the attribute reference and the other as the identity reference. The remaining frames are treated as targets to be generated, conditioned on a keypoint sequence representing the motion of the video. To prevent identity leakage, we apply complementary masking to the two reference frames so that the network must disentangle and combine the attribute and identity information to reconstruct the original video.
To enable this framework, we design a Dual ReferenceNet architecture that explicitly encodes the attribute and portrait references through two separate branches and fuses their disentangled features for generation via spatial attention. This design enables the network to move beyond simple pose driving, generating keypoint-driven portrait animations that seamlessly combine the attribute from one image with the identity from the other. Surprisingly, although the model is trained with only a single attribute reference at a time, the spatial attention mechanism allows more advanced operations at inference time. Since different attributes (e.g., hair, glasses, beard, hats) occupy distinct spatial regions, their features can be jointly injected without conflict, enabling seamless multi-attribute transfer. Furthermore, by interpolating the features of two attribute references, our model can achieve attribute interpolation, generating smooth transitions between the attributes. These emergent capabilities make our framework especially valuable for real-world styling scenarios, where users may want to explore diverse combinations and gradual transformations of facial attributes.
While self-reconstruction training is effective for learning to separate identity and attributes, it operates within a single video, leading to a domain gap when the model is applied to cross-identity inference, where the attribute and portrait come from different individuals. To mitigate this gap, we introduce a mask expansion strategy and lightweight augmentation schemes. These techniques expose the model to a broader range of attribute configurations during training, enabling robust transfer across spatial and structural variations of the attribute region. These designs form a unified framework capable of robust cross-identity attribute transfer. As a result, our method achieves a versatile system that generates portrait animations with diverse appearance edits in a zero-shot manner.
We summarize the key contributions of our work, as follows: (1) we propose the first method to generate keypoint-driven portrait animations with transferred attributes directly from two images, generalized across diverse facial attributes beyond hair; (2) we design a Dual ReferenceNet architecture that disentangles attribute and identity through two branches fused via spatial attention, enabling self-reconstruction training directly on uncurated in-the-wild videos without paired data; (3) we propose a mask expansion strategy and lightweight augmentations to bridge the domain gap for cross-identity transfer, improving robustness to diverse spatial configurations; and (4) our framework exhibits an emergent ability to support multi-attribute composition and interpolation in a single generation pass, without requiring any additional training.
@@ 2 Related Work
Face Editing.
@@
Diffusion-based Attribute Transfer.
Diffusion-based attribute transfer methods typically formulate editing as masked inpainting, where reference content is inserted into a target image using explicit masks (Yang et al., 2023; Chen et al., 2024; Mou et al., 2025; Chen et al., 2025; Song et al., 2025). These approaches have been adapted to domain-specific tasks such as hairstyle (Zhang et al., 2025; Chung et al., 2025), clothing (Kim et al., 2024a; Li et al., 2024; Chong et al., 2024), and makeup (Zhang et al., 2024b). While effective for static images, they rely on category labels or mask annotations. Video extensions (Fang et al., 2024; Tu et al., 2025) apply per-frame inpainting with post-hoc smoothing, but predefined masks are hard to specify for deformable facial attributes that vary over time. Recent works have also explored attribute transfer in 3D avatars (Kim et al., 2024b; Nam et al., 2025; Cha et al., 2024; 2025; Wang et al., 2025; Kim et al., 2025), but such approaches often require specialized capture setups or are not easily generalizable to in-the-wild scenarios. In contrast, our model performs attribute transfer and animation jointly in a single forward pass, conditioned only on a pair of reference images and a facial keypoint sequence. This eliminates the need for per-frame masks, text prompts, or category labels, enabling zero-shot transfer of diverse facial attributes.
@@
Portrait animation aims to generate motion from a static image, typically guided by facial keypoints, audio, or motion trajectories. Early methods rely on GANs with implicit keypoint modeling (Guo et al., 2024; Wang et al., 2021), while recent approaches use diffusion models (Hu, 2024; Zhu et al., 2024; Yang et al., 2025) for improved realism and temporal stability. These methods primarily focus on reenactment and identity preservation. Others incorporate paired motion (Xie et al., 2024) or audio (Yang et al., 2025), but require multi-stage inference or fine-tuning. Our model jointly performs facial attribute transfer and motion generation, producing photorealistic, identity-preserving videos from diverse attribute references and keypoint-driven motion in a single pass.
@@ 3 Method
Figure 2: Overview of Training Pipeline. Given an attribute-masked portrait image $\tilde{\mathbf{I}}_{\mathrm{port}}$ and an attribute-only image $\tilde{\mathbf{I}}_{\mathrm{attr}}$, Durian synthesizes a portrait animation with the transferred attribute. These inputs are constructed by randomly sampling two frames from a training video and applying the estimated masks. A sequence of facial keypoints $\{{\bm{k}_{\tau}}\}_{\tau=1}^{F}$ is extracted from the video to guide the motion. During generation, spatial features from PRNet and ARNet are fused via spatial attention into the DNet, ensuring identity preservation and attribute consistency in the synthesized video.
3.1 Overview: Learning Attribute Transfer from Self-Reconstruction
We propose a diffusion-based generative framework for portrait animation with cross-identity attribute transfer. At a high level, our model generates an $F$-frame animation sequence $\mathbf{V}=\{\mathbf{I}_{\tau}\}_{\tau=1}^{F}$ as:
@@ 3.1 Overview: Learning Attribute Transfer from Self-Reconstruction
conditioned on an attribute image $\mathbf{I}_{\mathrm{attr}}$, a portrait image $\mathbf{I}_{\mathrm{port}}$, and a sequence of driving facial keypoint images $\mathbf{K}=\{\bm{k}_{\tau}\}_{\tau=1}^{F}$. Each reference image has a binary mask: $\mathbf{M}_{\mathrm{attr}}$ localizes the attribute region (e.g., hair or glasses) in the reference image, while $\mathbf{M}_{\mathrm{port}}$ specifies the candidate region in the portrait where the attribute will be transferred. Using these masks, we construct two masked inputs: the attribute-only image $\tilde{\mathbf{I}}_{\mathrm{attr}}=\mathbf{I}_{\mathrm{attr}}\odot\mathbf{M}_{\mathrm{attr}}$, where only the attribute region is preserved, and the attribute-masked portrait image $\tilde{\mathbf{I}}_{\mathrm{port}}=\mathbf{I}_{\mathrm{port}}\odot(1-\mathbf{M}_{\mathrm{port}})$, where the corresponding region is removed. These masked inputs are fed into the Dual ReferenceNet, consisting of the Attribute ReferenceNet (ARNet) and Portrait ReferenceNet (PRNet), which extract multi-scale spatial features. These features are then injected into a diffusion-based generator, the Denoising UNet (DNet), to synthesize the remaining frames of the video with keypoint guidance $\mathbf{K}$ (Section˜3.2).
@@
To enable training without requiring explicitly annotated triplets (i.e., combinations of a target attribute image, an original portrait image, and an edited portrait image), we adopt a self-reconstruction strategy based on portrait videos (Yu et al., 2023; Xie et al., 2022). Specifically, we simulate attribute transfer by sampling two frames $\mathbf{I}_{\mathrm{attr}}$ and $\mathbf{I}_{\mathrm{port}}$ from the same video, treating one as the attribute reference and the other as the target portrait. We then construct the masked inputs $\tilde{\mathbf{I}}_{\mathrm{attr}}$ and $\tilde{\mathbf{I}}_{\mathrm{port}}$ using the same masking formulation as in inference, based on a segmentation mask of a randomly selected attribute. Although the two frames come from the same identity, the complementary masking enforces a clear separation between identity and attribute inputs, encouraging the model to learn meaningful mappings from these features to output frames without requiring cross-identity supervision. To enhance the model’s ability to generalize beyond the self-attribute transfer setup, we introduce an augmentation scheme that improves robustness to spatial and appearance variations(Section˜3.3).
@@
At inference time, we estimate refined attribute masks by aligning the attribute image to the portrait through a lightweight alignment process, mitigating spatial misalignment between them. Conditioned on the two masked reference images and the driving keypoint sequence, our model then synthesizes portrait animations with attribute transfer. Notably, our design also supports multi-attribute composition and smooth interpolation within a single generation pass, without requiring additional training or post-processing (Section˜3.4). Fig.˜1 shows our generated portrait animations with attribute transfer.
@@ 3.2 Model Architecture: Dual ReferenceNet
Inspired by recent approaches (Guo et al., 2023; Hu, 2024; Zhu et al., 2024) that leverage ReferenceNet to inject spatial features into diffusion models, we propose a Dual ReferenceNet architecture tailored for portrait animation with attribute transfer. Unlike previous work, our model includes two separate encoders: Attribute ReferenceNet (ARNet) and Portrait ReferenceNet (PRNet), each sharing the same architecture as the Denoising U-Net (DNet) in the diffusion model, excluding the temporal layers. The networks follow the U-Net (Long et al., 2015) architecture used in latent diffusion models (Rombach et al., 2022), with each block containing convolutional layers followed by self- and cross-attention modules. The overall architecture is shown in Fig.˜2.
@@
Given an attribute image $\mathbf{I}_{\mathrm{attr}}\in\mathbb{R}^{3\times H\times W}$ and a portrait image $\mathbf{I}_{\mathrm{port}}\in\mathbb{R}^{3\times H\times W}$, along with their binary masks $\mathbf{M}_{\mathrm{attr}}\in\mathbb{R}^{1\times H\times W}$ and $\mathbf{M}_{\mathrm{port}}\in\mathbb{R}^{1\times H\times W}$, which localize the attribute region and the candidate transfer region respectively, we construct two masked inputs: the attribute-only image $\tilde{\mathbf{I}}_{\mathrm{attr}}=\mathbf{I}_{\mathrm{attr}}\odot\mathbf{M}_{\mathrm{attr}}$, where only the attribute region is preserved, and the attribute-masked portrait image $\tilde{\mathbf{I}}_{\mathrm{port}}=\mathbf{I}_{\mathrm{port}}\odot(1-\mathbf{M}_{\mathrm{port}})$, where the corresponding candidate region is removed. We then encode these masked images into latent representations using the pretrained VAE from the latent diffusion model (Rombach et al., 2022), yielding $\bm{z}_{\mathrm{attr}},\bm{z}_{\mathrm{port}}\in\mathbb{R}^{c\times h\times w}$. The corresponding masks $\mathbf{M}_{\mathrm{attr}},\mathbf{M}_{\mathrm{port}}$ are downsampled to match the latent resolution, producing $\bm{m}_{\mathrm{attr}},\bm{m}_{\mathrm{port}}\in\mathbb{R}^{1\times h\times w}$. These downsampled masks are concatenated with the latents along the channel dimension to form $(c+1)$-channel inputs $\tilde{\bm{z}}_{\mathrm{attr}},\tilde{\bm{z}}_{\mathrm{port}}\in\mathbb{R}^{(c+1)\times h\times w}$ as follows:
$\tilde{\bm{z}}_{\mathrm{attr}}=\mathrm{concat}_{\mathrm{c}}(\bm{z}_{\mathrm{attr}},\bm{m}_{\mathrm{attr}}),\quad\tilde{\bm{z}}_{\mathrm{port}}=\mathrm{concat}_{\mathrm{c}}(\bm{z}_{\mathrm{port}},\bm{m}_{\mathrm{port}}).$ (2)
@@
The augmented latents are passed to ARNet $\mathcal{E}_{\mathrm{attr}}$ and PRNet $\mathcal{E}_{\mathrm{port}}$ to extract multi-scale feature maps after convolutional layers of each block:
$\mathcal{F}_{\mathrm{attr}}\coloneq\{\mathbf{F}_{\mathrm{attr}}^{l}\}_{l=1}^{L}=\mathcal{E}_{\mathrm{attr}}(\tilde{z}_{\mathrm{attr}};\Theta_{\mathrm{attr}}),\quad\mathcal{F}_{\mathrm{port}}\coloneq\{\mathbf{F}_{\mathrm{port}}^{l}\}_{l=1}^{L}=\mathcal{E}_{\mathrm{port}}(\tilde{z}_{\mathrm{port}};\Theta_{\mathrm{port}}),$ (3)
where $\Theta_{\{\mathrm{attr,port}\}}$ are the parameters of Dual ReferenceNet. Let $\mathbf{F}_{t}^{\tau,l}\in\mathbb{R}^{c_{l}\times h_{l}\times w_{l}}$ denote the feature map of the frame $\tau$ at the $l$-th block of the denoising U-Net. While the original denoising U-Net includes a self-attention layer at each resolution, we replace it with our spatial attention to integrate identity and attribute features in a spatially-aware manner. We denote width-wise concatenation as $\mathrm{concat}_{\mathrm{w}}(\cdot)$, and define our spatial attention $\mathrm{SA}(\cdot,\cdot,\cdot)$ as:
@@
$\bar{\mathbf{F}}_{t}^{\tau,l}=\mathrm{SA}(\mathbf{F}_{t}^{\tau,l},\mathbf{F}_{\mathrm{port}}^{l},\mathbf{F}_{\mathrm{attr}}^{l})=\mathrm{Attention}(\mathbf{W}_{Q}\mathbf{F}_{t}^{\tau,l},\mathbf{W}_{K}\mathbf{F}_{\mathrm{ref},t}^{\tau,l},\mathbf{W}_{V}\mathbf{F}_{\mathrm{ref},t}^{\tau,l}),$ (5)
where $\bar{\mathbf{F}}_{t}^{\tau,l}\in\mathbb{R}^{c_{l}\times h_{l}\times w_{l}}$ is the feature map after the spatial attention, $\mathrm{Attention}(Q,K,V)=\mathrm{softmax}({QK^{\top}}/{\sqrt{d}})V$ is the standard scaled dot-product attention (Vaswani et al., 2017), $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ are linear projection layers. This width-wise concatenation preserves spatial resolution and allows the model to attend across all positions in the combined reference and target features. As a result, the model can leverage both attribute and portrait guidance at every step.
@@
After applying spatial attention, we further inject semantic guidance into both the Dual ReferenceNet and the denoising U-Net via cross-attention. For ARNet, we use the CLIP (Radford et al., 2021) embedding of the attribute-only image $\tilde{\mathbf{I}}_{\mathrm{attr}}$ as the attribute embedding $\bm{\phi}_{\mathrm{attr}}$, which is injected via cross-attention into each block of ARNet. For PRNet and DNet, we construct a portrait embedding $\bm{\phi}_{\mathrm{port}}$ by combining ArcFace (Deng et al., 2019) and CLIP embeddings of the attribute-masked portrait image $\tilde{\mathbf{I}}_{\mathrm{port}}$ following StableAnimator (Tu et al., 2024). This embedding is injected into both PRNet and DNet to enhance identity preservation. We define the cross-attention operation $\mathrm{CA}(\cdot,\cdot)$ as: